OpenStudy (anonymous):

|5x-3|=12 inequality solutions would be x=3 and x=-9/5, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah that is correct

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

-12=5x-3=12 -9=5x=15 -9/5=x x=3